OpenStudy (anitas.):

At a meeting of the local computer user group, each member brought two non-members. If a total of 27 people attended, how many were members and how many were non-members?

OpenStudy (anitas.):

@inkyvoyd

OpenStudy (notshh):

let x = # of members 2x = # of nonmembers Set equal to 27 so x + 2x = 27 make sense or should I continue?

OpenStudy (anitas.):

i need more info on how i find my answer with this prob and the steps

OpenStudy (notshh):

ok then simplify x + 2x = 27 to 3x = 27 solve for x by dividing both sides by 3 x = 9

OpenStudy (anitas.):

ok so i divide it and i get x=9 so now i know that there is 9 members but how do i find the non-members? @notshh

OpenStudy (anitas.):

oh duh Actually i know. you have to subtract 27-9=18 so I have 9 members and 18 non-members.. Am i right?

OpenStudy (notshh):

correct. (: alternatively, you could do 9 x 2 since x = # of members and 2x = # of nonmembers, but subtraction works just as well.

OpenStudy (anitas.):

ok thank you sooooo much for helping me!!! <3333333

OpenStudy (notshh):

no problem!